Due Date: Midnight Sunday, October 27th Assume you are in the market for a brand new Toyota Tacoma Limited pickup truck, because you’ve heard that only really cool people drive them. You’ve sat down with a sales man at your local Toyota dealership and he has offered you three different options for the purchase of this vehicle. The price you’ve agreed on is $35,000. You will be putting 20% down. Assume there are no prepayment penalties. 1) What is the principal amount of the loan you will be taking out? i.e. How much will you be borrowing? •

1.PV =

2) Option 1 is a 48 month loan with an interest rate of 4.57% annually. What would be your monthly payment? How much interest will you pay over the life of the loan? • PMT = • Interest paid =

3) Option 2 is a 60 month loan with an interest rate of 4.61% annually. What would be your monthly payment? How much interest will you pay over the life of the loan • PMT = • Interest paid =

4) Option 3 is a 72 month loan with an interest rate of 4.70% annually. What would be your monthly payment? How much interest will you pay over the life of the loan? • PMT = • Interest paid =

5) Use the provided amortization schedule to amortize the 60 month loan (question 2) – however, assume you pay $100 of additional principal every month. How many months will it take you to pay off the loan? How much interest will you save?

6) Can you think of a reason why it might make sense to take the longer-term loan, as opposed to the shorter-term loan, even though you know you’ll effectively pay more interest?
Answer #1

1.
=35000*(1-20%)=28000

2.
PMT=35000*(1-20%)*(4.57%/12)/(1-1/(1+4.57%/12)^48)=639.380457240767

Interest=35000*(1-20%)*(4.57%/12)/(1-1/(1+4.57%/12)^48)*48-35000*(1-20%)=2690.26194755682

3.
PMT=35000*(1-20%)*(4.61%/12)/(1-1/(1+4.61%/12)^60)=523.406216972053

Interest=35000*(1-20%)*(4.61%/12)/(1-1/(1+4.61%/12)^60)*60-35000*(1-20%)=3404.37301832321

4.
PMT=35000*(1-20%)*(4.7%/12)/(1-1/(1+4.7%/12)^72)=447.052043575293

Interest=35000*(1-20%)*(4.7%/12)/(1-1/(1+4.7%/12)^72)*72-35000*(1-20%)=4187.74713742111

P.S.: I am not allowed to answer more than 4 questions